Find the standard form of the equation of the ellipse with the given characteristics.
vertces: (-3, -10), (-3,0) minor axis of length 4
Please help with this problem.

Answer by lwsshak3(11628)

Find the standard form of the equation of the ellipse with the given characteristics.
vertces: (-3, -10), (-3,0) minor axis of length 4
Please help with this problem.
..
Given information shows this is an ellipse with center at (-3,-5), and its major axis is vertical.
Standard form of the equation for this ellipse: (x-h)^2/a^2+(y-k)^2/b^2=1 (a>b), where (h,k) are the (x,y) coordinates of the center.
length of major axis=10=2a
a=5
a^2=25
length of minor axis=4=2b
b=2
b^2=4
c^2=a^2-b^2=25-4=21
c=sqrt(21)=4.58..(length of foci)
Equation:
(x+3)^2/25+(y+5)^2/4=1
See graph of this ellipse below:
..
y=+-(4(1-(x+3)^2/25))^.5-5